# MemBench: Memorized Image Trigger Prompt Dataset for Diffusion Models

### v1

@@ Abstract
Diffusion models have achieved remarkable success in Text-to-Image generation tasks, leading to the development of many commercial models. However, recent studies have reported that diffusion models often generate replicated images in train data when triggered by specific prompts, potentially raising social issues ranging from copyright to privacy concerns. To sidestep the memorization, there have been recent studies for developing memorization mitigation methods for diffusion models. Nevertheless, the lack of benchmarks impedes the assessment of the true effectiveness of these methods. In this work, we present MemBench, the first benchmark for evaluating image memorization mitigation methods. Our benchmark includes a large number of memorized image trigger prompts in Stable Diffusion, the most popularly used model nowadays. Furthermore, in contrast to the prior work evaluating mitigation performance only on trigger prompts, we present metrics evaluating on both trigger prompts and general prompts, so that we can see whether mitigation methods address the memorization issue while maintaining performance for general prompts. This is an important development considering the practical applications which previous works have overlooked. Through evaluation on MemBench, we verify that the performance of existing image memorization mitigation methods is still insufficient for application to diffusion models.
@@ 1 Introduction
Text-to-Image (T2I) generation has shown significant advancements and successes with the advance of diffusion models, such as Stable Diffusion [18]. Compared to previous generative models, diffusion models excel in generating diverse and high quality images from user-desired text prompts, which has led to the vast release of commercial models such as MidJourney. However, recent studies [20, 21, 1] have revealed that certain text prompts tend to keep replicating images in the train dataset which can cause serious privacy issues. This issue has already triggered controversy in the real world: specific prompts containing the term “Afghan” have been known to reproduce copyrighted images of the Afghan girl when using MidJourney [29]. One of the major issues with such prompts is that, regardless of initial random noise leveraged in the reverse process of the diffusion model, they always invoke almost or exactly same memorized images [29, 1, 26].
@@
To address this matter, Wen et al. [29] and Somepalli et al. [21] have proposed mitigation methods to prevent the regeneration of identical images in the train dataset invoked from certain text prompts. However, the evaluation of these memorization mitigation methods has lacked rigor and comprehensiveness due to the absence of benchmarks. As an adhoc assessment method, the current studies [29, 21] have adopted the following method: 1) simulating memorization by fine-tuning T2I diffusion models for overfitting on a separate small and specific dataset of {image, prompt} pairs, and 2) assess whether the images used in the fine-tuning are reproduced from the query prompts. However, it remains unclear whether such evaluation results can be extended to practical scenarios with the pre-trained existing diffusion models and represent the effectiveness for resolving memorization.
@@
In this work, we present MemBench, the first benchmark for evaluating image memorization mitigation methods for diffusion models. Our MemBench includes the following key features to ensure effective evaluation: (1) MemBench provides a sufficient number of memorized image trigger prompts, with 3000 and 1500 prompts for Stable Diffusion 1 and 2, most popular open-source models, respectively. Previous work [26] has offered the highest number of prompts at 345 and 218 for Stable Diffusion 1 and 2, respectively; however, we increase the number of prompts to enhance the reliability of the evaluation. (2) We take into account a general prompt scenario, which has been overlooked in prior work. The prior mitigation methods [29, 16, 21] have been evaluated solely on memorized image trigger prompts. However, to apply mitigation methods in practice, they should also maintain performance on general prompts. (3) We suggest rigorous metrics. As previous mitigation works [29, 21] have measured, MemBench also includes the SSCD [14], which measures the similarity between memorized and generated images, and the CLIP Score [7], which measures Text-Image alignment. Additionally, MemBench involves Aesthetic Score [19] to assess image quality, which has been overlooked by prior work and allows to penalize unuseful trivial solutions. (4) We propose the reference performance that mitigation methods should achieve. In previous works [16, 29, 21], the effectiveness of mitigation methods has been demonstrated by measuring the decrease in SSCD and the extent to which the CLIP Score is maintained before and after applying the mitigation method. However, this does not necessarily confirm whether image memorization has been adequately mitigated. Therefore, for SSCD and CLIP Score, we provide guidelines on the target value that should be achieved for a mitigation method to be considered effective.
@@
Our additional contribution lies in offering an effective algorithm to search for memorized image trigger prompts. The absence of such benchmarks originates from the significant challenge of collecting prompts that induce memorized images. Compared to the existing methods [1, 27] that require training data itself [19] for fetching candidate prompts, our proposed based construction method is developed with Markov Chain Monte Carlo (MCMC) [12], which enables the efficient search for these problematic prompts in an open token space without any dataset.
@@ 2 Related Work
Memorization Mitigation Methods. Somepalli et al. [21] propose a rule-based text embedding augmentation to mitigate memorization in diffusion models. This includes adding Gaussian noise to text embeddings or inserting random tokens in the prompt. Wen et al. [29] propose the loss that predicts if a prompt will induce a memorized image, and present a mitigation strategy that applies adversarial attacks on this loss to modify the text embeddings of trigger prompts. Both of these works evaluate their methods by intentionally overfitting the diffusion model on specific {image, text} pairs to induce the memorization effect, and then checking whether the images are regenerated from the corresponding prompts when their methods are applied. Ren et al. [16] analyze the impact of trigger prompts on the cross-attention layer of diffusion models and propose a corresponding mitigation method.
@@
Training Data Extraction Attack. Carlini et al. [1] propose a method to search for memorized image trigger prompts involving two stages. In the pre-processing stage, they embed the entire training set of Stable Diffusion into the CLIP [15] feature space and cluster these embeddings to identify the most repeated images. In the post-processing stage, Stable Diffusion is used to generate 500 images for each prompt corresponding to these clustered images. The similarity among these 500 generated images is measured, and only those prompts that produce highly similar images are sampled. Finally, image retrieval is performed on the training data using generated images from these selected prompts to verify if the generated images match the training data images. The pre-processing involves CLIP embedding and clustering of 160M images, while the post-processing involves generating 175M images. Webster [26] propose an advanced searching algorithm. In the pre-processing stage, an encoder is trained to compress CLIP embeddings. Then, 2B CLIP embeddings are compressed and clustered using KNN [27]. In the post-processing stage, Webster introduces an effective method that performs a few inferences of the diffusion model to predict whether a prompt will induce memorized images. This method is applied to 20M prompts acquired from the pre-processing stage.
[…]
@@
Benchmark Dataset. Since the only existing dataset that can be used for evaluating mitigation methods is the small dataset released by Webster [26], Ren et al. [16] evaluate their method on the Webster dataset, while the Webster dataset is not originally purposed as a benchmark dataset. The dataset is constructed by the training data extraction attack method proposed by Webster, which is not scalable; thus, the dataset remains a small scale. Also, Ren et al. did not measure the loss of semantic preservation after mitigation, which is an important criterion but overlooked. Our benchmark is also constructed by a novel and efficient data extraction method, and the first benchmark for evaluating those mitigation methods with carefully designed metrics.
@@ 3.1 Preliminary
Denoising Diffusion Probabilistic Model (DDPM) [9] is a representative diffusion model designed to approximate the real data distribution $q({\mathbf{x}})$ with a model $p_{\bm{\theta}}({\mathbf{x}})$. For each ${\mathbf{x}}_{0}\sim q({\mathbf{x}})$, DDPM constructs a discrete Markov chain $\{{\mathbf{x}}_{0},{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{T}\}$ that satisfies $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{t-1},\beta_{t}{\mathbf{I}})$. This is referred to as the forward process, where $\{\beta_{t}\}_{t=1}^{T}$ is a sequence of positive noise scales. Conversely, the reverse process generates images according to $p_{\bm{\theta}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})=\mathcal{N}({\mathbf{x}}_{t};\mu_{\bm{\theta}}({\mathbf{x}}_{t},t),\Sigma_{\bm{\theta}}({\mathbf{x}}_{t},t))$. DDPM starts by sampling ${\mathbf{x}}_{T}$ from a Gaussian distribution, and then undergoes a stochastic reverse process to generate the sample ${\mathbf{x}}_{0}$, i.e. an image. With a parametrized denoising network ${\bm{\epsilon}}_{{\bm{\theta}}}$, this generation process can be expressed as:
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{t}$, $\sigma_{t}$ can be $\sqrt{\beta}$ or $\sqrt{\frac{1-\bar{\alpha_{t-1}}}{1-\bar{\alpha}_{t}}\beta_{t}}$, and ${\mathbf{w}}\sim\mathcal{N}(0;{\mathbf{I}})$. The equations may vary depending on hyper-parameter choices and the numerical solver used [22, 23].
@@
In T2I diffusion models such as Stable Diffusion [18], CFG [8] is commonly employed to generate images better aligned with the desired prompt. Given a text prompt ${\mathbf{p}}$ and the text encoder ${\mathbf{f}}(\cdot)$ of the pre-trained CLIP [15], predicted noise is replaced as follows:
@@
Wen et al. [29] proposed an efficient method to predict whether a prompt will generate an image included in the training data. Prior to presenting this method, we present the definition of image memorization suggested in [21, 1].
@@
Given a train set $\mathcal{D}_{train}=\{({\mathbf{x}}_{train,i},{\mathbf{p}}_{train,i})\}_{i=1}^{N}$, a generated image ${\mathbf{x}}$ from a diffusion model ${\bm{\epsilon}}_{\bm{\theta}}$ trained on $\mathcal{D}_{train}$, and a similarity measurement score SSCD [14], image memorization of ${\mathbf{x}}$ is defined as:
@@
The prior works [1, 29, 26] found that prompts inducing memorized images do so regardless of the initial noise, ${\mathbf{x}}_{T}$, i.e., repeatedly generating the same or almost identical images despite different ${\mathbf{x}}_{T}$. To quickly identify this case, Wen et al. [29] propose a measure to predict whether a prompt will induce a memorized image using only the first step of the diffusion model, without generating the image. This measure, referred to as $D_{\bm{\theta}}$, is formulated as follows:
@@ 3.2 Memorization Trigger Prompt Searching as an Optimization Problem
[…]
@@ 3.3 Constructing MCMC Leveraging $D_{\theta}$
To tackle the aforementioned challenging optimization problem, we propose to use Markov Chain Monte Carlo (MCMC) [6] to sample from the target distribution $\pi({\mathbf{p}})$. This method allows us to efficiently explore the discrete prompt space and find prompts likely to induce memorized images, effectively navigating $\mathcal{P}$ to identify optimal prompts. From any arbitrary distribution of sentence, $\pi_{0}$, Markov Chain with transition matrix ${\mathbf{T}}$ can be developed as follows:
@@
It is well known that Markov Chains satisfying irreducibility and aperiodicity converge to certain distribution $\pi^{*}$ [17], which can be formulated as $\pi_{n}=\pi_{0}{\mathbf{T}}^{n}\rightarrow\pi^{*}$ independent of $\pi_{0}$. In this context, the transition matrix can vary depending on the algorithm used to solve the MCMC. By carefully choosing the sampling algorithm, we can ensure that the final distribution $\pi^{*}$ reached by the transition matrix converges to desired target distribution $\pi$ [17, 4, 6]. Considering the multi-dimensional nature of our parameter space, we employ the Gibbs sampling algorithm [4] for simplicity. Gibbs sampling is an MCMC sampling algorithm method where, at each step, only one coordinate of the multi-dimensional variable is updated to transition from the current state to the next state. Gibbs sampling algorithm has proven the convergence of the transition matrix and is known for fast convergence in multi-dimensional problems [10, 24, 13]. We adopt random scan Gibbs sampling, which involves randomly selecting an index and updating the value at that index. This process can be expressed as the sum of $n$ transition matrices, as follows:
@@
$[{\mathbf{T}}_{i}]_{{\mathbf{p}}^{j}\rightarrow{\mathbf{p}}^{j+1}}$ =$=\left\{\begin{array}[]{cl}\pi({\mathbf{p}}^{j+1}|{\mathbf{p}}^{j})&\text{if }{\mathbf{p}}_{-i}^{j}={\mathbf{p}}^{j+1}_{-i}\\ 0&\text{else},\end{array}\right.$ (10)
@@
where detailed derivation is provided in Supplementary Material. Since it is impractical to compute $D_{\bm{\theta}}(\cdot)$ for all ${\mathbf{w}}\in\mathcal{W}$, we approximate $\mathcal{W}$ as top $Q$ samples obtained from BERT [2]. This means that the $i$-th element of the prompt ${\mathbf{p}}$ is masked and BERT is used to predict the word, from which the top Q samples are selected as candidate words. Mathematical derivation is complex, but the algorithm is straightforward: the process iteratively 1) selects and replace a word into [MASK] token from the sentence, 2) predicts top $Q$ words via BERT and computes proposal distribution, and 3) stochastically replaces it according to the proposal distribution. Please refer to Algorithm 1 for details.

### v2

@@ Abstract
Diffusion models have achieved remarkable success in Text-to-Image generation tasks, leading to the development of many commercial models. However, recent studies have reported that diffusion models often generate replicated images in train data when triggered by specific prompts, potentially raising social issues ranging from copyright to privacy concerns. To sidestep the memorization, there have been recent studies for developing memorization mitigation methods for diffusion models. Nevertheless, the lack of benchmarks impedes the assessment of the true effectiveness of these methods. In this work, we present MemBench, the first benchmark for evaluating image memorization mitigation methods. Our benchmark includes a large number of memorized image trigger prompts in various Text-to-Image diffusion models. Furthermore, in contrast to the prior work evaluating mitigation performance only on trigger prompts, we present metrics evaluating on both trigger prompts and general prompts, so that we can see whether mitigation methods address the memorization issue while maintaining performance for general prompts. This is an important development considering the practical applications which previous works have overlooked. Through evaluation on MemBench, we verify that the performance of existing image memorization mitigation methods is still insufficient for application to diffusion models. The code and datasets are available at https://github.com/chunsanHong/MemBench_code
$\dagger$\$$\dagger$\$footnotetext: Co-corresponding authors.
@@ 1 Introduction
Text-to-Image (T2I) generation has shown significant advancements and successes with the advance of diffusion models, such as Stable Diffusion (Rombach et al., 2022). Compared to previous generative models, diffusion models excel in generating diverse and high quality images from user-desired text prompts, which has led to the vast release of commercial models such as MidJourney. However, recent studies (Somepalli et al., 2023a; b; Carlini et al., 2023) have revealed that certain text prompts tend to keep replicating images in the train dataset which can cause serious privacy issues. This issue has already triggered controversy in the real world: specific prompts containing the term “Afghan” have been known to reproduce copyrighted images of the Afghan girl when using MidJourney (Wen et al., 2024b). One of the major issues with such prompts is that, regardless of initial random noise leveraged in the reverse process of the diffusion model, they always invoke almost or exactly same memorized images (Wen et al., 2024b; Carlini et al., 2023; Webster, 2023).
@@
To address this matter, Wen et al. (2024b) and Somepalli et al. (2023b) have proposed mitigation methods to prevent the regeneration of identical images in the train dataset invoked from certain text prompts. However, the evaluation of these memorization mitigation methods has lacked rigor and comprehensiveness due to the absence of benchmarks. As an adhoc assessment method, the current studies (Wen et al., 2024b; Somepalli et al., 2023b) have adopted the following method: 1) simulating memorization by fine-tuning T2I diffusion models for overfitting on a separate small and specific dataset of {image, prompt} pairs, and 2) assess whether the images used in the fine-tuning are reproduced from the query prompts. However, it remains unclear whether such evaluation results can be extended to practical scenarios with the pre-trained existing diffusion models and represent the effectiveness for resolving memorization.
@@
In this work, we present MemBench, the first benchmark for evaluating image memorization mitigation methods for diffusion models. Our MemBench includes the following key features to ensure effective evaluation: (1) MemBench provides a sufficient number of memorized image trigger prompts, with 3000 and 1500 prompts for Stable Diffusion 1 and 2, most popular open-source models, respectively. Previous work (Webster, 2023) has offered the highest number of prompts at 345 and 218 for Stable Diffusion 1 and 2, respectively; however, we increase the number of prompts to enhance the reliability of the evaluation. We also provide memorized images and trigger prompts for DeepFloydIF (Shonenkov et al., 2023) and Realistic Vision (CivitAI, 2023). (2) We take into account a general prompt scenario, which has been overlooked in prior work. The prior mitigation methods (Wen et al., 2024b; Ren et al., 2024; Somepalli et al., 2023b) have been evaluated solely on memorized image trigger prompts. However, to apply mitigation methods in practice, they should also maintain performance on general prompts. (3) We suggest rigorous metrics. As previous mitigation works (Wen et al., 2024b; Somepalli et al., 2023b) have measured, MemBench also includes the SSCD (Pizzi et al., 2022), which measures the similarity between memorized and generated images, and the CLIP Score (Hessel et al., 2021), which measures Text-Image alignment. Additionally, MemBench involves Aesthetic Score (Schuhmann et al., 2022) to assess image quality, which has been overlooked by prior work and allows to penalize unuseful trivial solutions. (4) We propose the reference performance that mitigation methods should achieve. In previous works (Ren et al., 2024; Wen et al., 2024b; Somepalli et al., 2023b), the effectiveness of mitigation methods has been demonstrated by measuring the decrease in SSCD and the extent to which the CLIP Score is maintained before and after applying the mitigation method. However, this does not necessarily confirm whether image memorization has been adequately mitigated. Therefore, for SSCD and CLIP Score, we provide guidelines on the target value that should be achieved for a mitigation method to be considered effective.
@@
Our additional contribution lies in offering an effective algorithm to search for memorized image trigger prompts. The absence of such benchmarks originates from the significant challenge of collecting prompts that induce memorized images. Compared to the existing methods (Carlini et al., 2023; Webster et al., 2023) that require training data itself (Schuhmann et al., 2022) for fetching candidate prompts, our proposed based construction method is developed with Markov Chain Monte Carlo (MCMC) (Metropolis et al., 1953), which enables the efficient search for these problematic prompts in an open token space without any dataset.
@@ 2 Related Work
Memorization Mitigation Methods. Memorization mitigation methods are divided into two approaches: one is the inference time method, and the other is the train time method. Inference time methods aim to prevent the generation of memorized images that already exist in pretrained diffusion models during the image generation process. Somepalli et al. (2023b) propose a rule-based text embedding augmentation to mitigate memorization in diffusion models. This includes adding Gaussian noise to text embeddings or inserting random tokens in the prompt. Wen et al. (2024b) propose the loss that predicts if a prompt will induce a memorized image, and present a mitigation strategy that applies adversarial attacks on this loss to modify the text embeddings of trigger prompts. Both of these works evaluate their methods by intentionally overfitting the diffusion model on specific {image, text} pairs to induce the memorization effect, and then checking whether the images are regenerated from the corresponding prompts when their methods are applied. Ren et al. (2024) analyze the impact of trigger prompts on the cross-attention layer of diffusion models and propose a corresponding mitigation method.
Train time methods aim to prevent diffusion models from memorizing training data during model training by employing specific training techniques. Although several methods (Daras et al., 2024; Liu et al., 2024) have been proposed, experiments have been conducted only on small models and datasets such as CIFAR-10 and CelebHQ. While some experiments (Somepalli et al., 2023b; Ren et al., 2024; Wen et al., 2024b) have been conducted on large models such as Stable Diffusion, they only assess whether the fine-tuning dataset is memorized when fine-tuning the model. To date, no train time mitigation method has been tested by training large diffusion models from scratch to evaluate its effectiveness.
Therefore, we consider inference time methods as the primary approaches applicable to foundation models like Stable Diffusion. However, due to the absence of a proper benchmark that contains sufficient number of memorized image trigger prompts, evaluation remains inadequate. MemBench addresses this gap by providing sufficient test data and appropriate metrics for the comprehensive evaluation of these inference time methods.
@@
Training Data Extraction Attack. Carlini et al. (2023) propose a method to search for memorized image trigger prompts involving two stages. In the pre-processing stage, they embed the entire training set of Stable Diffusion into the CLIP (Radford et al., 2021) feature space and cluster these embeddings to identify the most repeated images. In the post-processing stage, Stable Diffusion is used to generate 500 images for each prompt corresponding to these clustered images. The similarity among these 500 generated images is measured, and only those prompts that produce highly similar images are sampled. Finally, image retrieval is performed on the training data using generated images from these selected prompts to verify if the generated images match the training data images. The pre-processing involves CLIP embedding and clustering of 160M images, while the post-processing involves generating 175M images. Webster (2023) propose an advanced searching algorithm. In the pre-processing stage, an encoder is trained to compress CLIP embeddings. Then, 2B CLIP embeddings are compressed and clustered using KNN (Webster et al., 2023). In the post-processing stage, Webster introduces an effective method that performs a few inferences of the diffusion model to predict whether a prompt will induce memorized images. This method is applied to 20M prompts acquired from the pre-processing stage.
Both methods share common bottlenecks: they are memory inefficient and require extremely high computational costs. Additionally, the most fundamental problem is their reliance on training data as candidate trigger prompts. With LAION becoming inaccessible [^1] https://laion.ai/notes/laion-maintenance/, these methods can no longer be reproducible and utilized. However, our method can search more for trigger prompts efficiently than those methods even without any pre-processing steps and any dataset.
@@
Benchmark Dataset. Since the only existing dataset that can be used for evaluating mitigation methods is the small dataset released by Webster (2023), Ren et al. (2024) evaluate their method on the Webster dataset, while the Webster dataset is not originally purposed as a benchmark dataset. The dataset is constructed by the training data extraction attack method proposed by Webster, which is not scalable; thus, the dataset remains a small scale. Also, Ren et al. did not measure the loss of semantic preservation after mitigation, which is an important criterion but overlooked. Our benchmark is also constructed by a novel and efficient data extraction method, and the first benchmark for evaluating those mitigation methods with carefully designed metrics.
@@ 3.1 Preliminary
Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) is a representative diffusion model designed to approximate the real data distribution $q({\mathbf{x}})$ with a model $p_{\bm{\theta}}({\mathbf{x}})$. For each ${\mathbf{x}}_{0}\sim q({\mathbf{x}})$, DDPM constructs a discrete Markov chain $\{{\mathbf{x}}_{0},{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{T}\}$ that satisfies $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{t-1},\beta_{t}{\mathbf{I}})$. This is referred to as the forward process, where $\{\beta_{t}\}_{t=1}^{T}$ is a sequence of positive noise scales. Conversely, the reverse process generates images according to $p_{\bm{\theta}}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})=\mathcal{N}({\mathbf{x}}_{t};\mu_{\bm{\theta}}({\mathbf{x}}_{t},t),\Sigma_{\bm{\theta}}({\mathbf{x}}_{t},t))$. DDPM starts by sampling ${\mathbf{x}}_{T}$ from a Gaussian distribution, and then undergoes a stochastic reverse process to generate the sample ${\mathbf{x}}_{0}$, i.e. an image. With a parametrized denoising network ${\bm{\epsilon}}_{{\bm{\theta}}}$, this generation process can be expressed as:
@@
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{t}$, $\sigma_{t}$ can be $\sqrt{\beta}$ or $\sqrt{\frac{1-\bar{\alpha_{t-1}}}{1-\bar{\alpha}_{t}}\beta_{t}}$, and ${\mathbf{w}}\sim\mathcal{N}(0;{\mathbf{I}})$. The equations may vary depending on hyper-parameter choices and the numerical solver used (Song et al., 2021a; b).
@@
In T2I diffusion models such as Stable Diffusion (Rombach et al., 2022), CFG (Ho & Salimans, 2022) is commonly employed to generate images better aligned with the desired prompt. Given a text prompt ${\mathbf{p}}$ and the text encoder ${\mathbf{f}}(\cdot)$ of the pre-trained CLIP (Radford et al., 2021), predicted noise is replaced as follows:
@@
Wen et al. (2024b) proposed an efficient method to predict whether a prompt will generate an image included in the training data. Prior to presenting this method, we present the definition of image memorization suggested in (Somepalli et al., 2023b; Carlini et al., 2023).
@@
Given a train set $\mathcal{D}_{train}=\{({\mathbf{x}}_{train,i},{\mathbf{p}}_{train,i})\}_{i=1}^{N}$, a generated image ${\mathbf{x}}$ from a diffusion model ${\bm{\epsilon}}_{\bm{\theta}}$ trained on $\mathcal{D}_{train}$, and a similarity measurement score SSCD (Pizzi et al., 2022), image memorization of ${\mathbf{x}}$ is defined as:
@@
The prior works (Carlini et al., 2023; Wen et al., 2024b; Webster, 2023) found that prompts inducing memorized images do so regardless of the initial noise, ${\mathbf{x}}_{T}$, i.e., repeatedly generating the same or almost identical images despite different ${\mathbf{x}}_{T}$. To quickly identify this case, Wen et al. (2024b) propose a measure to predict whether a prompt will induce a memorized image using only the first step of the diffusion model, without generating the image. This measure, referred to as $D_{\bm{\theta}}$, is formulated as follows:
@@ 3.2 Memorization Trigger Prompt Searching as an Optimization Problem
Our objective is to construct a memorized image trigger prompts dataset and verify corresponding memorized images, i.e. to construct $\mathcal{D}_{mem}=\{{\mathbf{p}}\mid\mathbb{E}[\mathcal{M}({\mathbf{x}}({\bm{\epsilon}}_{\bm{\theta}},{\mathbf{p}}),\mathcal{D}_{train})]>\kappa,{\mathbf{p}}\in\mathcal{T}\}$ where $\kappa$ is the threshold and $\mathcal{T}$ is space of all possible prompts. As mentioned in Section 2, the prior works (Carlini et al., 2023; Webster, 2023) utilized $\mathcal{D}_{train}$ to search for candidate prompts that could become $\mathcal{D}_{mem}$. They then generated images from these candidate prompts and conducted image retrieval to find memorized images within $\mathcal{D}_{train}$ which is expensive. Moreover, since the training dataset, LAION, is no longer accessible, this approach becomes infeasible. Thus, we approach the problem from a different perspective. We search for candidate prompts that could become $\mathcal{D}_{mem}$ without using $\mathcal{D}_{train}$. Then, we generate images from these candidate prompts and use a Reverse Image Search API [^2] https://tineye.com/ to find images on the web akin to generated ones by regarding the web as the training set. Finally, we perform a human verification process.
@@ 3.3 Constructing MCMC Leveraging $D_{\theta}$
To tackle the aforementioned challenging optimization problem, we propose to use Markov Chain Monte Carlo (MCMC) (Hastings, 1970) to sample from the target distribution $\pi({\mathbf{p}})$. This method allows us to efficiently explore the discrete prompt space and find prompts likely to induce memorized images, effectively navigating $\mathcal{P}$ to identify optimal prompts. From any arbitrary distribution of sentence, $\pi_{0}$, Markov Chain with transition matrix ${\mathbf{T}}$ can be developed as follows:
@@
It is well known that Markov Chains satisfying irreducibility and aperiodicity converge to certain distribution $\pi^{*}$ (Robert et al., 1999), which can be formulated as $\pi_{n}=\pi_{0}{\mathbf{T}}^{n}\rightarrow\pi^{*}$ independent of $\pi_{0}$. In this context, the transition matrix can vary depending on the algorithm used to solve the MCMC. By carefully choosing the sampling algorithm, we can ensure that the final distribution $\pi^{*}$ reached by the transition matrix converges to desired target distribution $\pi$ (Robert et al., 1999; Geman & Geman, 1984; Hastings, 1970). Considering the multi-dimensional nature of our parameter space, we employ the Gibbs sampling algorithm (Geman & Geman, 1984) for simplicity. Gibbs sampling is an MCMC sampling algorithm method where, at each step, only one coordinate of the multi-dimensional variable is updated to transition from the current state to the next state. Gibbs sampling algorithm has proven the convergence of the transition matrix and is known for fast convergence in multi-dimensional problems (Johnson et al., 2013; Terenin et al., 2020; Papaspiliopoulos & Roberts, 2008). We adopt random scan Gibbs sampling, which involves randomly selecting an index and updating the value at that index. This process can be expressed as the sum of $n$ transition matrices, as follows:
@@
$[{\mathbf{T}}_{i}]_{{\mathbf{p}}^{j}\rightarrow{\mathbf{p}}^{j+1}}$ =$=\left\{\begin{array}[]{cl}\pi({\mathbf{p}}^{j+1}_{i}|{\mathbf{p}}^{j}_{-i})&\text{if }{\mathbf{p}}_{-i}^{j}={\mathbf{p}}^{j+1}_{-i}\\ 0&\text{else},\end{array}\right.$ (10)
@@
where detailed derivation is provided in Appendix. Since it is impractical to compute $D_{\bm{\theta}}(\cdot)$ for all ${\mathbf{w}}\in\mathcal{W}$, we approximate $\mathcal{W}$ as top $Q$ samples obtained from BERT (Devlin et al., 2018). This means that the $i$-th element of the prompt ${\mathbf{p}}$ is masked and BERT is used to predict the word, from which the top Q samples are selected as candidate words. Mathematical derivation is complex, but the algorithm is straightforward: the process iteratively 1) selects and replace a word into [MASK] token from the sentence, 2) predicts top $Q$ words via BERT and computes proposal distribution, and 3) stochastically replaces it according to the proposal distribution. Please refer to Algorithm 1 for details.